# Modelling Cosmic-Ray Transport: Magnetized versus Unmagnetized Motion in Astrophysical Magnetic Turbulence

### v2

@@ Abstract
Cosmic-ray transport in turbulent astrophysical environments remains a multifaceted problem, and despite decades of study the impact of complex magnetic field geometry — evident in simulations and observations — has only recently received more focused attention. To understand how ensemble-averaged transport behaviour emerges from the intricate interactions between cosmic rays and structured magnetic turbulence, we run test-particle experiments in snapshots of a strongly turbulent magnetohydrodynamics simulation. We characterize particle-turbulence interactions via the gyro radii of particles and their experienced field-line curvatures, which reveals two distinct transport modes: magnetized motion, where particles are tightly bound to strong coherent flux tubes and undergo large-scale mirroring; and unmagnetized motion characterized by chaotic scattering through weak and highly tangled regions of the magnetic field. We formulate an effective stochastic process for each mode: compound subdiffusion with long mean free paths for magnetized motion, and a Langevin process with short mean free paths for unmagnetized motion. A combined stochastic walker that alternates between these two modes accurately reproduces the mean squared displacements observed in the test-particle data. Our results emphasize the critical role of coherent magnetic structures in comprehensively understanding cosmic-ray transport and lay a foundation for developing a theory of geometry-mediated transport.
@@ 1 Introduction
Turbulence inevitably appears wherever non-linear flows are at play, which applies to virtually all astrophysical plasmas, such as the solar wind (Engelbrecht et al., 2022), the interstellar medium (Elmegreen & Scalo, 2004), or the intracluster medium (Ryu et al., 2012). It plays a key role in answering questions about the physics of cosmic rays, such as those regarding transport and energization processes, as well as the effects of self-generated turbulence and dynamically relevant pressure (Amato & Blasi, 2018). Applications, including Hopkins et al. (2020); Dörner et al. (2023); Ewart et al. (2024); and Kamal Youssef & Grenier (2024), rely on a solid understanding of the transport of fast charged particles through turbulent magnetic fields, which has been intensively studied since Parker (1965). However, due to the highly complex nature of plasma turbulence and the strong dependence of cosmic-ray transport on turbulence properties, a comprehensive description is difficult to attain. This is illustrated by the circumstance that, despite extensive work, existing phenomenological theories are hardly compatible with the available observed cosmic-ray data, as recently discussed by Kempski & Quataert (2022) and Hopkins et al. (2022).
@@
In large parts of the existing literature, the principal paradigm in understanding the interplay between cosmic rays and magnetic turbulence has been gyro-resonance, i.e., scattering of particles by Alfvén waves for comparable gyro radii and wavenumbers (Kulsrud, 2005). These waves can either be self-generated by streaming cosmic rays (Skilling, 1975), or emerge as the constituting parts of an extrinsic turbulent cascade. The latter case is classically treated with quasi-linear theory (QLT), where small-angle scattering in pitch angle caused by gyro-resonance accumulates to a random walk of the particle along a magnetic field line (Jokipii, 1966; Mertsch, 2020; Reichherzer et al., 2020; Els et al., 2024). However, this paradigm views magnetic turbulence merely as an ensemble of linear space-filling magnetohydrodynamic (MHD) waves with a given energy spectrum. This is incompatible with strong MHD turbulence (Meyrand et al., 2016), which tends to form coherent structures, i.e., ordered patches characterized by strong alignment and reduced non-linearity (see, e.g., Perez & Boldyrev, 2009; Matthaeus et al., 2015). A geometric perspective on magnetic turbulence is also advocated by, e.g., Grauer et al. (1994); Politano & Pouquet (1995); Grauer & Marliani (2000); Mininni et al. (2006); Boldyrev (2006); and Malara et al. (2021). Coherent structures have been extensively observed in the solar wind (Tu & Marsch, 1995; Khabarova et al., 2021; Vinogradov et al., 2024), and numerically studied in the context of cosmic-ray acceleration (Arzner et al., 2006; Lemoine, 2021; Pezzi et al., 2022; Pugliese et al., 2023).
@@
While strong turbulence as a mixture of coherent and chaotic structures marginally reproduces the expected energy spectrum, such an averaging argument does not readily apply to cosmic-ray transport, because the transport coefficients depend strongly on the geometry and distribution of these structures. This was demonstrated by Shukurov et al. (2017), who found longer mean free paths (MFP) of test particles in a fluctuating dynamo compared to random-phase synthetic turbulence with the same energy spectrum. Additionally, coherent structures may act as non-resonant magnetic mirrors (Chandran et al., 1999; Albright et al., 2001; Bell et al., 2025) and govern field-line wandering, which are two important mechanisms observed in test-particle experiments in snapshots of MHD simulations (Beresnyak et al., 2011; Xu & Yan, 2013; Cohet & Marcowith, 2016; Zhang & Xu, 2024). The wandering of field lines is also important for the effective diffusion of streaming cosmic rays (Sampson et al., 2023). Further, sharply curved magnetic field lines, which may arise at the interfaces between coherent structures, were recently invoked as an effective scattering agent for cross-field transport (Kempski et al., 2023; Lemoine, 2023). Even though a micro-physical theory of cosmic-ray transport beyond gyro-resonance is not yet available, the general understanding of magnetic turbulence has improved significantly in recent years (see Schekochihin 2022 for a comprehensive review), thus providing the foundations for the development of such a theory. That is not to say that gyro-resonance is entirely obsolete; rather, actual transport behaviour likely results from the interplay of several different mechanisms with varying contributions, depending on the turbulent properties of a given astrophysical system, which may vary in space and scale.
@@
In this paper, we explore how a theory of cosmic-ray transport mediated by turbulence-induced geometry may look like. We do this by means of test-particle experiments in an MHD simulation of a fluctuating dynamo, where a dynamically dominant flow amplifies an initially small magnetic field and produces pronounced coherent flux tubes (Schekochihin et al., 2004; Rincon, 2019; Seta et al., 2020). This process is believed to play an important role in the generation of magnetic fields in galaxies (Rieder & Teyssier, 2017; Gent et al., 2024) and the intracluster medium (Vazza et al., 2018; Steinwandel et al., 2024), with initial seed fields possibly provided by the Weibel instability (Sironi et al., 2023; Zhou et al., 2024). To understand the connection between test-particle motion and magnetic field geometry, we investigate magnetic moment variations and mean squared displacements (MSD) conditional on the gyro radii of the particles and their experienced field-line curvature. This reveals two distinct modes of transport, namely magnetized motion, where particles are closely bound to a strong and ordered magnetic field inside coherent structures, and unmagnetized motion consisting of chaotic scattering through a weak and tangled magnetic field. This description can be made more precise by fitting a stochastic model to the test-particle motion, where field-line wandering and mirroring in the magnetized case are represented by compound subdiffusion (Balescu et al., 1994; Qin et al., 2002; Neuer & Spatschek, 2006; Minnie et al., 2009), and the chaotic scattering in the unmagnetized case is represented by a 3D Langevin equation (Chandrasekhar, 1943; Bian & Li, 2024). We conclude our treatment by discussing implications and open questions towards a proper theory for cosmic-ray transport.
@@ 2 MHD and test-particle Simulations
2.1 MHD Simulation Setup
We perform a simulation of incompressible visco-resistive MHD turbulence in a 3D periodic box. The governing equations of the flow field $\boldsymbol{u}$ and the magnetic field $\boldsymbol{B}$, which are given by (Biskamp, 2003)
@@ 2.1 MHD Simulation Setup
with hyper-viscosity $\nu_{h}$ and hyper-resistivity $\eta_{h}$ of order $h$, are solved with our pseudo-spectral code SpecDyn (Wilbert et al., 2022; Wilbert, 2023). The equations are solved on a 3D uniform grid with $1024^{3}$ points and periodic boundary conditions. The length of the domain is $L_{\mathrm{box}}=2\pi$ and the time scale of the flow is $T_{\mathrm{eddy}}=L_{\mathrm{box}}/u_{\mathrm{rms}}$. We set $\nu_{h}=\eta_{h}$, resulting in the magnetic Prandtl number $Pr_{m}=1$. The magnetic field $\boldsymbol{B}$, which is specified in Alfvénic units $[B]=[u]$, is initialized with a small-amplitude seed field with zero magnetic helicity and zero net magnetic flux. The flow field $\boldsymbol{u}$ is driven by a large-scale force density $\boldsymbol{f}$, operating on the wavenumber band $1\leq k\leq 2$, with random amplitudes, which are $\delta$-correlated in time to ensure constant power injection. The force density is constructed such that the injected net cross-helicity is zero (Alvelius, 1999). Due to the chosen wavenumber band of the forcing, the box contains only one flow correlation cell. We run the simulations until the magnetic energy has been amplified to a statistically saturated state, and then record 8 snapshots of $\boldsymbol{u}(\boldsymbol{x})$ and $\boldsymbol{B}(\boldsymbol{x})$ separated in time by $T_{\mathrm{eddy}}$.
We simulate equations (1) with $h=1$ as the baseline case and with $h=2$ as the production case. The parameters of the simulations are listed in table 1. Setting $h=2$ results in a sharper dissipative cut-off and, thus, an extended inertial range, which is reflected by the Reynolds numbers listed in table 1. Differences in the spectra and geometry are further reflected by the characteristic scales listed in table 2. For $h=2$ the magnetic energy saturates at a higher level compared to $h=1$ (see also Brandenburg & Sarson, 2002), however this does not matter for our test particle simulations, because we normalize the magnetic field to unit root mean square strength. Despite differences in the flux rope geometries as indicated in table 2, we do not observe significant differences in our conditional particle statistics. Thus, subsequently we only show the results for the case $h=2$, which additionally allows for the meaningful inclusion of lower-energy particles due to the shorter dissipation scale.
@@
Table 2: Characteristic length scales of the simulations. We report the correlation scale $k_{w,\mathrm{corr}}$, Taylor scale $k_{w,T}$ and Kolmogorov dissipation scale $k_{w,\mathrm{diss}}$ for both fields $w\in\{u,B\}$. Additionally, we show for the magnetic field the characteristic parallel scale $k_{\parallel}$, reversal scale $k_{\boldsymbol{B}\times\boldsymbol{j}}$ and perpendicular scale $k_{\boldsymbol{B}\cdot\boldsymbol{j}}$ (see Schekochihin et al., 2004).
[…]
2.2 Test Particle Setup
with the volume-preserving Boris scheme (Boris & Shanny, 1971; Ripperda et al., 2018). This test-particle approach is justified by assuming that relativistic cosmic rays with $V\approx c$ move in front of a non-relativistic plasma background with $u_{\mathrm{rms}}\ll c$, such that the time scales of the MHD and test particle simulations are well separated, i.e. $T_{\mathrm{particle}}\ll T_{\mathrm{eddy}}$ with $T_{\mathrm{particle}}=L_{\mathrm{box}}/V_{0}$. We can then neglect the influence of the electric field, implying conservation of energy for our test particles. The particles are parametrized by the normalized root mean square (RMS) gyro frequency
which includes the amplitudes $V_{0}$ of the particle velocity and $B_{\mathrm{rms}}$ of the magnetic field; equation (2) is accordingly normalized to $\|\boldsymbol{V}\|=1$ and $B_{\mathrm{rms}}=\langle\|\boldsymbol{B}\|^{2}\rangle^{1/2}=1$. The parameter $\hat{\omega}_{g}$ denotes how strongly particles are coupled to the magnetic field and inversely encodes their energy via $E=\gamma mc^{2}=qB_{\mathrm{rms}}L_{\mathrm{box}}c/\hat{\omega}_{g}$, with $V_{0}=c$. Table 3 provides a mapping to specific astrophysical systems, based on typical values of the magnetic field strengths and turbulence correlation scales. Due to limited numerical resolution, and the requirement to constrain the particle gyro radii to the range of resolved turbulent fluctuations, the representable particle energies are very high for the respective physical systems. How our findings extend to lower energies, such as galactic GeV protons, can only be speculated about, as we attempt in Section 4.1.
[…]
@@ 2.2 Test Particle Setup
Table 3: Cosmic-ray protons affected by our considerations with $\hat{\omega}_{g}=64,\cdots,256$, characterized by the RMS field strength $B_{\mathrm{rms}}$ and turbulence correlation scale $L_{\mathrm{corr}}\sim L_{\mathrm{box}}$ with values taken from (1) Weygand et al. (2011), (2) Weygand et al. (2013), (3) Houde et al. (2009), (4) Crutcher (2012), (5) Jansson & Farrar (2012), (6) Seta et al. (2020), and (7) Dom´ınguez Fernández (2020). Note that values found in the literature tend to be widely spread due to the heterogeneous nature of astrophysical environments, as well as due to intrinsic observational uncertainties. The listed values serve merely as order-of-magnitude estimates. Also note that our fluctuating dynamo turbulence is not applicable to the anisotropic solar wind (Chen et al., 2020; Wang et al., 2024), and that the reported particle energies are non-relativistic with $V_{0}=(0.037,\cdots,0.009)\,c$. For these reasons, the solar wind values are shown only for reference.
We solve equation (2) with our ParticlePusher code, which is able to record binned statistics of arbitrary quantities along particle trajectories. We utilize the local gyro average of a quantity $Q$ along a particle trajectory $\boldsymbol{X}_{t}$, which is defined as

### v3

# Modelling cosmic-ray transport: magnetised versus unmagnetised motion in astrophysical magnetic turbulence

@@ Abstract
Cosmic-ray transport in turbulent astrophysical environments remains a multifaceted problem and, despite decades of study, the impact of complex magnetic field geometry — evident in simulations and observations — has only recently received more focussed attention. To understand how ensemble-averaged transport behaviour emerges from the intricate interactions between cosmic rays and structured magnetic turbulence, we run test-particle experiments in snapshots of a strongly turbulent magnetohydrodynamics simulation. We characterise particle–turbulence interactions via the gyro radii of particles and their experienced field-line curvatures, which reveals two distinct transport modes: magnetised motion, where particles are tightly bound to strong coherent flux tubes and undergo large-scale mirroring; and unmagnetised motion, characterised by chaotic scattering through weak and highly tangled regions of the magnetic field. We formulate an effective stochastic process for each mode: compound subdiffusion with long mean free paths for magnetised motion, and a Langevin process with short mean free paths for unmagnetised motion. A combined stochastic walker that alternates between these two modes accurately reproduces the mean squared displacements observed in the test-particle data. Our results emphasise the critical role of coherent magnetic structures in comprehensively understanding cosmic-ray transport and lay a foundation for developing a theory of geometry-mediated transport.
@@ 1 Introduction
Turbulence inevitably appears wherever nonlinear flows are at play, which applies to virtually all astrophysical plasmas, such as the solar wind (Engelbrecht et al., 2022), the interstellar medium (Elmegreen & Scalo, 2004), or the intracluster medium (Ryu et al., 2012). It plays a key role in answering questions about the physics of cosmic rays, such as those regarding transport and energisation processes, as well as the effects of self-generated turbulence and dynamically relevant pressure (Amato & Blasi, 2018). Applications, including those reported by Hopkins et al. (2020), Dörner et al. (2023), Ewart et al. (2024) and Kamal Youssef & Grenier (2024), rely on a solid understanding of the transport of fast charged particles through turbulent magnetic fields, which has been intensively studied since the work of Parker (1965). However, due to the highly complex nature of plasma turbulence and the strong dependence of cosmic-ray transport on turbulence properties, a comprehensive description is difficult to attain. This is illustrated by the circumstance that, despite extensive work, existing phenomenological theories are hardly compatible with the available observed cosmic-ray data, as recently discussed by Kempski & Quataert (2022) and Hopkins et al. (2022).
@@
In large parts of the existing literature, the principal paradigm in understanding the interplay between cosmic rays and magnetic turbulence has been gyro-resonance, i.e. scattering of particles by Alfvén waves for comparable gyro radii and wavenumbers (Kulsrud, 2005). These waves can either be self-generated by streaming cosmic rays (Skilling, 1975), or emerge as the constituting parts of an extrinsic turbulent cascade. The latter case is classically treated with quasi-linear theory (QLT), where small-angle scattering in pitch angle caused by gyro-resonance accumulates to a random walk of the particle along a magnetic field line (Jokipii, 1966; Mertsch, 2020; Reichherzer et al., 2020; Els et al., 2024). However, this paradigm views magnetic turbulence merely as an ensemble of linear space-filling magnetohydrodynamic (MHD) waves with a given energy spectrum. This is incompatible with strong MHD turbulence (Meyrand et al., 2016), which tends to form coherent structures, i.e. ordered patches characterised by strong alignment and reduced nonlinearity (see, e.g. Perez & Boldyrev, 2009; Matthaeus et al., 2015). A geometric perspective on magnetic turbulence is also advocated by, e.g. Grauer et al. (1994), Politano & Pouquet (1995), Grauer & Marliani (2000), Mininni et al. (2006), Boldyrev (2006) and Malara et al. (2021). Coherent structures have been extensively observed in the solar wind (Tu & Marsch, 1995; Khabarova et al., 2021; Vinogradov et al., 2024), and numerically studied in the context of cosmic-ray acceleration (Arzner et al., 2006; Lemoine, 2021; Pezzi et al., 2022; Pugliese et al., 2023).
@@
While strong turbulence as a mixture of coherent and chaotic structures marginally reproduces the expected energy spectrum, such an averaging argument does not readily apply to cosmic-ray transport, because the transport coefficients depend strongly on the geometry and distribution of these structures. This was demonstrated by Shukurov et al. (2017), who found longer mean free paths (MFPs) of test particles in a fluctuating dynamo compared with random-phase synthetic turbulence with the same energy spectrum. Additionally, coherent structures may act as non-resonant magnetic mirrors (Chandran et al., 1999; Albright et al., 2001; Bell et al., 2025) and govern field-line wandering, which are two important mechanisms observed in test-particle experiments in snapshots of MHD simulations (Beresnyak et al., 2011; Xu & Yan, 2013; Cohet & Marcowith, 2016; Zhang & Xu, 2024). The wandering of field lines is also important for the effective diffusion of streaming cosmic rays (Sampson et al., 2023). Further, sharply curved magnetic field lines, which may arise at the interfaces between coherent structures, were recently invoked as an effective scattering agent for cross-field transport (Kempski et al., 2023; Lemoine, 2023). Even though a micro-physical theory of cosmic-ray transport beyond gyro-resonance is not yet available, the general understanding of magnetic turbulence has improved significantly in recent years (see Schekochihin 2022 for a comprehensive review), thus providing the foundations for the development of such a theory. That is not to say that gyro-resonance is entirely obsolete; rather, actual transport behaviour likely results from the interplay of several different mechanisms with varying contributions, depending on the turbulent properties of a given astrophysical system, which may vary in space and scale.
@@
In this paper, we explore how a theory of cosmic-ray transport mediated by turbulence-induced geometry may look like. We do this by means of test-particle experiments in an MHD simulation of a fluctuating dynamo, where a dynamically dominant flow amplifies an initially small magnetic field and produces pronounced coherent flux tubes (Schekochihin et al., 2004; Rincon, 2019; Seta et al., 2020). This process is believed to play an important role in the generation of magnetic fields in galaxies (Rieder & Teyssier, 2017; Gent et al., 2024) and the intracluster medium (Vazza et al., 2018; Steinwandel et al., 2024), with initial seed fields possibly provided by the Weibel instability (Sironi et al., 2023; Zhou et al., 2024). To understand the connection between test-particle motion and magnetic field geometry, we investigate magnetic moment variations and mean squared displacements (MSDs) conditional on the gyro radii of the particles and their experienced field-line curvature. This reveals two distinct modes of transport, namely magnetised motion, where particles are closely bound to a strong and ordered magnetic field inside coherent structures, and unmagnetised motion consisting of chaotic scattering through a weak and tangled magnetic field. This description can be made more precise by fitting a stochastic model to the test-particle motion, where field-line wandering and mirroring in the magnetised case are represented by compound subdiffusion (Balescu et al., 1994; Qin et al., 2002; Neuer & Spatschek, 2006; Minnie et al., 2009), and the chaotic scattering in the unmagnetised case is represented by a three-dimensional (3-D) Langevin equation (Chandrasekhar, 1943; Bian & Li, 2024). We conclude our treatment by discussing implications and open questions towards a proper theory for cosmic-ray transport.
@@ 2 MHD and test-particle simulations
2.1 MHD simulation set-up
We perform a simulation of incompressible visco-resistive MHD turbulence in a 3-D periodic box. The governing equations of the flow field $\boldsymbol{u}$ and the magnetic field $\boldsymbol{B}$, which are given by (Biskamp, 2003)
@@ 2.1 MHD simulation set-up
with hyper-viscosity $\nu_{h}$ and hyper-resistivity $\eta_{h}$ of order $h$, are solved with our pseudo-spectral code SpecDyn (Wilbert et al., 2022; Wilbert, 2023). The equations are solved on a 3-D uniform grid with $1024^{3}$ points and periodic boundary conditions. The length of the domain is $L_{\mathrm{box}}=2\pi$ and the time scale of the flow is $T_{\mathrm{eddy}}=L_{\mathrm{box}}/u_{\mathrm{rms}}$. We set $\nu_{h}=\eta_{h}$, resulting in the magnetic Prandtl number $Pr_{m}=1$. The magnetic field $\boldsymbol{B}$, which is specified in Alfvénic units $[B]=[u]$, is initialised with a small-amplitude seed field with zero magnetic helicity and zero net magnetic flux. The flow field $\boldsymbol{u}$ is driven by a large-scale force density $\boldsymbol{f}$, operating on the wavenumber band $1\leq k\leq 2$, with random amplitudes, which are $\delta$-correlated in time to ensure constant power injection. The force density is constructed such that the injected net cross-helicity is zero (Alvelius, 1999). Due to the chosen wavenumber band of the forcing, the box contains only one flow correlation cell. We run the simulations until the magnetic energy has been amplified to a statistically saturated state, and then record eight snapshots of $\boldsymbol{u}(\boldsymbol{x})$ and $\boldsymbol{B}(\boldsymbol{x})$ separated in time by $T_{\mathrm{eddy}}$.
We simulate equations (1) with $h=1$ as the baseline case and with $h=2$ as the production case. The parameters of the simulations are listed in table 1. Setting $h=2$ results in a sharper dissipative cut-off and, thus, an extended inertial range, which is reflected by the Reynolds numbers listed in table 1. Differences in the spectra and geometry are further reflected by the characteristic scales listed in table 2. For $h=2$ the magnetic energy saturates at a higher level compared with $h=1$ (see also Brandenburg & Sarson, 2002); however, this does not matter for our test particle simulations, because we normalise the magnetic field to unit root-mean-square (r.m.s.) strength. Despite differences in the flux rope geometries, as indicated in table 2, we do not observe significant differences in our conditional particle statistics. Thus, subsequently, we only show the results for the case $h=2$, which additionally allows for the meaningful inclusion of lower-energy particles due to the shorter dissipation scale.
@@
Table 2: Characteristic length scales of the simulations. We report the correlation scale $k_{w,\mathrm{corr}}$, Taylor scale $k_{w,T}$ and Kolmogorov dissipation scale $k_{w,\mathrm{diss}}$ for both fields $w\in\{u,B\}$. Additionally, we show for the magnetic field, the characteristic parallel scale $k_{\parallel}$, reversal scale $k_{\boldsymbol{B}\times\boldsymbol{j}}$ and perpendicular scale $k_{\boldsymbol{B}\cdot\boldsymbol{j}}$ (see Schekochihin et al., 2004).
@@
Table 3: Cosmic-ray protons affected by our considerations with $\hat{\omega}_{g}=64,\cdots,256$, characterised by the r.m.s. field strength $B_{\mathrm{rms}}$ and turbulence correlation scale $L_{\mathrm{corr}}\sim L_{\mathrm{box}}$ with values taken from: (1) Weygand et al. (2011); (2) Weygand et al. (2013); (3) Houde et al. (2009); (4) Crutcher (2012); (5) Jansson & Farrar (2012); (6) Seta et al. (2020); and (7) Domínguez Fernández (2020). Note that values found in the literature tend to be widely spread due to the heterogeneous nature of astrophysical environments, as well as due to intrinsic observational uncertainties. The listed values serve merely as order-of-magnitude estimates. Also note that our fluctuating dynamo turbulence is not applicable to the anisotropic solar wind (Chen et al., 2020; Wang et al., 2024), and that the reported particle energies are non-relativistic with $V_{0}=(0.037,\cdots,0.009)\,c$. For these reasons, the solar wind values are shown only for reference.
2.2 Test particle set-up
with the volume-preserving Boris scheme (Boris & Shanny, 1971; Ripperda et al., 2018). This test-particle approach is justified by assuming that relativistic cosmic rays with $V\approx c$ move in front of a non-relativistic plasma background with $u_{\mathrm{rms}}\ll c$, such that the time scales of the MHD and test particle simulations are well separated, i.e. $T_{\mathrm{particle}}\ll T_{\mathrm{eddy}}$ with $T_{\mathrm{particle}}=L_{\mathrm{box}}/V_{0}$. We can then neglect the influence of the electric field, implying conservation of energy for our test particles. The particles are parametrised by the normalised r.m.s. gyro frequency
which includes the amplitudes $V_{0}$ of the particle velocity and $B_{\mathrm{rms}}$ of the magnetic field; equation (2) is accordingly normalised to $\|\boldsymbol{V}\|=1$ and $B_{\mathrm{rms}}=\langle\|\boldsymbol{B}\|^{2}\rangle^{1/2}=1$. The parameter $\hat{\omega}_{g}$ denotes how strongly particles are coupled to the magnetic field and inversely encodes their energy via $E=\gamma mc^{2}=qB_{\mathrm{rms}}L_{\mathrm{box}}c/\hat{\omega}_{g}$, with $V_{0}=c$. Table 3 provides a mapping to specific astrophysical systems, based on typical values of the magnetic field strengths and turbulence correlation scales. Due to limited numerical resolution and the requirement to constrain the particle gyro radii to the range of resolved turbulent fluctuations, the representable particle energies are very high for the respective physical systems. How our findings extend to lower energies, such as galactic GeV protons, can only be speculated about, as we attempt in Section 4.1.
[…]
We solve equation (2) with our ParticlePusher code, which is able to record binned statistics of arbitrary quantities along particle trajectories. We use the local gyro average of a quantity $Q$ along a particle trajectory $\boldsymbol{X}_{t}$, which is defined as